Y2knic2:

Jake tossed a paper cup 50 times and recorded how it landed. The table shows the results: Position Open Side Up-3 Closed Side Up-7 Landing-40 Based on the table, determine the experimental probability of each outcome (landing open side up, landing closed side up, and landing on its side). Show your work. HELP PLSS

Y2knic2:

btw landing on its side is 40

Chloeisdum:

divid each number by 50

Chloeisdum:

3 divided by 50 is ?

Y2knic2:

0.6 right

Chloeisdum:

Yes, now (for example) turn each number into a precent 3/50=6/100= 6%

Chloeisdum:

7 divided by 50= .14

Chloeisdum:

0.06 for the first one

Y2knic2:

ty

Chloeisdum:

40 divided by 50 = .8

Y2knic2:

so all i had to do was find the decimals?

Chloeisdum:

Yes!

Y2knic2:

oh ok that makes sense thanks for the help

Chloeisdum:

Yw